OpenStudy (anonymous):

the product of (a-b)(a-b) is a perfect square trinomial, true, sometimes true, or never true?

OpenStudy (anonymous):

how come its trinomial? i think its binomial

OpenStudy (anonymous):

never true

OpenStudy (anonymous):

ok thanks, and i dont know either kanwar, its a question on a test

OpenStudy (anonymous):

NEVER TRUE

OpenStudy (anonymous):

hence, never true since its binomial

OpenStudy (anonymous):

wait, thats not right, i got it wrong, its not "never"

OpenStudy (anonymous):

ughh :(

OpenStudy (anonymous):

(a-b)(a-b)=(a-b)^2, so how come it is never a perfect square trinomial? I think always true.

OpenStudy (anonymous):

or like this (a-b)(a-b) = (a-b)^2 = a^2 - 2ab + b^2 so its true!!! my bad sorry!

OpenStudy (anonymous):

It is always true. Use FOIL: (a-b)*(a-b)= a*a+a*(-b) +(-b)*a+(-b)*(-b)= a^2-b*a-b*a+b^2=a^2-2ab+b^2 which will always factor as (a-b)(a-b).

OpenStudy (anonymous):

never true